The engine of a motor car, developing 7.6 kW, propels the car at 71 km/h along a level road against a head wind. If the car has a mass of 2.6 tonnes, what is the resistance to motion in Newtons per tonne?
1
Expert's answer
2020-07-22T10:34:31-0400

The engine power

"P=Fv"

Hence, the resistance force to motion

"F=\frac{P}{v}=\frac{7600\:\rm W}{71\times \frac{1000\:\rm m}{3600\:\rm s}}=385.35\:\rm N"

"F/m=\frac{385.35\:\rm N}{2.6\:\rm tonne}=148.2\:\rm N/tonne"